OpenStudy (anonymous):

A can of fruit contains 3 1/2 cups of fruit. The suggested serving size is 1/2 cup. How many servings are in the can of fruit? a 7 b 1 2/3 c 4 d 2 2/3

OpenStudy (noseboy908):

To start, you should make the 3 1/2 cups total into an improper fraction. Know how to do that?

OpenStudy (anonymous):

@mxskylar12 The answer is A. 3 1/2 / 1/2 = 7.

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} servings&ounces\\ \hline\\ 1&\frac{1}{2}\\ x&3\frac{1}{2} \end{array}\implies \cfrac{1}{x}=\cfrac{\frac{1}{2}}{3\frac{1}{2}}\)